A $110,000 investment is made. Over a 5-year period, a return of $30,000 occurs at the end of the first year. Each successive year yields a return that is 9% less than the previous year's return. If money is worth 5, what is the equivalent present worth for the investment? _______ 
Answer #1

PW ($) = - 110,000 + 30,000 x P/F(5%, 1) + 30,000 x 0.91% x P/F(5%, 2) + 30,000 x 0.91 x 0.91 x P/F(5%, 3) + 30,000 x 0.91 x 0.91 x 0.91 x P/F(5%, 4) + 30,000 x 0.91 x 0.91 x 0.91 x 0.91 x P/F(5%, 5)

= - 110,000 + 30,000 x 0.9524 + 27,300 x 0.9070 + 24,843 x 0.8638 + 22,607.13 x 0.8227 + 20,572.49 x 0.7835

= - 110,000 + 28,572 + 24,761.1 + 21,459.38 + 18,598.89 + 16,118.54

= - 490.09

= - 490